OpenStudy (anonymous):

A sector of a circle has a central angle of 135 degrees. Find the area of the sector if the radius of the circle is 11 cm.

hero (hero):

\[\frac{x}{360} = \frac{y}{\pi r^2}\] x = measure of sector = 135 y = area of sector

OpenStudy (anonymous):

\[135/360 * 11^2\pi \]

OpenStudy (anonymous):

because the area of the circle is pir^2

hero (hero):

@muiklover317, what you posted alone doesn't really mean much. It's important to show the ratios of proportionality involved in developing the formula.

OpenStudy (anonymous):

you don't necessarily need the proportion.

OpenStudy (anonymous):

that's the harder and more complicated way to see this problem.

OpenStudy (anonymous):

since it's 360 in circle,

hero (hero):

Actually it's easier and simpler.

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

it's different for everyone

OpenStudy (anonymous):

you can't say this is easier and that is harder. it depends on a person.

OpenStudy (anonymous):

that's your way and this is my way.

OpenStudy (anonymous):

jhalt is one that will choose the better way for him to solve.

hero (hero):

It's the same thing if you think about it. If we use your method, we'll have to remember different formulas for different things. If you use my formula, then you just remember the ratio of proportionality.

OpenStudy (anonymous):

it's not the formula and you don't need to memorize it. It's logical. You already know the area and circumference of the circle right?

OpenStudy (anonymous):

and you already know that it's gonna be 360 degrees

OpenStudy (anonymous):

for me, it's confusing to set the proportion cuz you can easily mess up with the order.

hero (hero):

No you can't.

OpenStudy (anonymous):

45.375 pi is the answer?

OpenStudy (anonymous):

I didn't get that @jhalt

OpenStudy (anonymous):

135 times pi times 121/360 = y

OpenStudy (anonymous):

@Hero Like I told you, that's YOUR way to solve it. Why do you even need to argue about this?

OpenStudy (anonymous):

oh nono @jhalt

OpenStudy (anonymous):

you see how 135 is the angle right?

OpenStudy (anonymous):

so you have to write (or combine) them together like this: 135/360 times 121pi

OpenStudy (anonymous):

cuz 121 is the area of the circle.

hero (hero):

Actually @jhalt has it

OpenStudy (anonymous):

so isn't it then 16335pi/360?

hero (hero):

Simplify that

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

it's around 142.55

OpenStudy (anonymous):

45.375 pi

OpenStudy (anonymous):

thanks for the help guys.

OpenStudy (anonymous):

I don't want to cause a dispute so I will not award the best response. Both methods were helpful and I appreciate them

OpenStudy (anonymous):

no prob :)

OpenStudy (anonymous):

oh you left the pi out okay

hero (hero):

The only thing is...@jhalt when you setup the proportion, what you want to do is simplify the fraction when possible. So if you set up this: \[\frac{135}{360} = \frac{y}{\pi r^2}\]

hero (hero):

Simplify the fraction on the left side first before isolating y

OpenStudy (anonymous):

ok so its smaller and easier right? But you don't have to correct?

hero (hero):

It simplifies to 3/8: \[\frac{3}{8} = \frac{y}{\pi r^2}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Both methods help

hero (hero):

@musiklover317, what if you were given the area of the sector but had to find the angle. Then what would you do?

OpenStudy (anonymous):

give me an example question

hero (hero):

The area of a sector of the circle is 10. The radius is 5. Find the angle of the sector.

hero (hero):

@musiklover317

OpenStudy (anonymous):

you don't have pi>?

OpenStudy (anonymous):

whateve

hero (hero):

It is exactly as you see it.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x = 144

OpenStudy (anonymous):

that's the answer okay?

hero (hero):

Is that supposed to be your answer to the question?

OpenStudy (anonymous):

yeah i got that

hero (hero):

That's not correct.

OpenStudy (anonymous):

i left out the pi

OpenStudy (anonymous):

look.

OpenStudy (anonymous):

x/360*25=10

hero (hero):

Bro, you don't know how to solve it.

hero (hero):

It's still a circle bro.

OpenStudy (anonymous):

then show it to me ha

OpenStudy (anonymous):

it's the area

hero (hero):

Just because I didn't include pi in my question doesn't mean you remove pi from the formula.

OpenStudy (anonymous):

oh then you'll have to divide by pi

OpenStudy (anonymous):

simple like that.

hero (hero):

Okay bro. You got it. But next time, don't remove pi from the formula just because a question doesn't have pi in it.

hero (hero):

Suppose you were given the angle of a sector and the area of a sector, but had to find the radius of the circle. Then what would you do?

hero (hero):

@musiklover317

hero (hero):

Suppose the angle is 216 degrees and the area of a sector is 15pi. What's the radius of the circle?